The angles are supplementary angles. Determine the measures of 1 and 2. m1=x m2=2x+15
Answer by cleomenius(959)

x + 2x + 15 = 180
3x = 165
x = 55 degrees. the smaller angle.
2(55) + 15 = 125 degrees the larger angle.
55 + 125 = 180 degrees.
Cleomenius.